OpenStudy (vishweshshrimali5):

Drawing direction field and solution curve passing through a given point

OpenStudy (vishweshshrimali5):

The question says: Graph a direction field (by a CAS or by hand). In the field graph several solution curves by hand, particularly those passing through the given points (x,y) . (1) \(y' = 1+y^2\)

OpenStudy (vishweshshrimali5):

Point is \((\cfrac{\pi}{4} , 1\))

OpenStudy (kainui):

Well in a sense, you have a formula for the slope right there! So you can just plug in that point: y'=1+(1)^2 The slope doesn't even depend on the x-value!

OpenStudy (vishweshshrimali5):

Now, I know that for drawing a direction field, the first step is to put 1+y^2 = 0 which will not give any solution. So, the arrows will never be horizontal. Now, I also know that 1+y^2 will always be positive, so, arrows will always point upwards. Now, I assume that 1+y^2 = 1 \(\implies y = 0\) Which means that arrows lying on x-axis will have a slope 1. Similarly now I assume 1+y^2 = 5 \(\implies y^2 = 4\) Which represents a pair of straight lines y = 2 and y = -2. So, arrows lying on these lines will have a slope of 5. And so on..

OpenStudy (vishweshshrimali5):

Now I don't know how to plot the solution curve.

OpenStudy (kainui):

Well can you plot the slope field?

OpenStudy (vishweshshrimali5):

Well sorry but I have to leave now. I will draw it and then notify you. Will that work ?

OpenStudy (kainui):

Yeah, sure. Are you going to eat dinner as well?

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (kainui):

I'll be back after I finish breakfast lol.

OpenStudy (vishweshshrimali5):

Lol Bye

OpenStudy (vishweshshrimali5):

This is the direction field.

OpenStudy (vishweshshrimali5):

The corresponding solution curve according to me.

OpenStudy (vishweshshrimali5):

@Kainui Mission Accomplished :)

OpenStudy (vishweshshrimali5):

Is that correct ?

OpenStudy (kainui):

Yeah, exactly! Not bad.

OpenStudy (vishweshshrimali5):

Thanks.

OpenStudy (kainui):

Glad I could help, but you did most of it yourself.

OpenStudy (vishweshshrimali5):

:)